# ECLARE: Efficient cross-planar learning for anisotropic resolution enhancement

### v1

@@ Abstract
In clinical imaging, magnetic resonance (MR) image volumes are often acquired as stacks of 2D slices, permitting decreased scan times, improved signal-to-noise ratio, and image contrasts unique to 2D MR pulse sequences. While this is sufficient for clinical evaluation, automated algorithms designed for 3D analysis perform sub-optimally on 2D-acquired scans, especially those with thick slices and gaps between slices. Super-resolution (SR) methods aim to address this problem, but previous methods do not address all of the following: slice profile shape estimation, slice gap, domain shift, and non-integer / arbitrary upsampling factors. In this paper, we propose ECLARE (Efficient Cross-planar Learning for Anisotropic Resolution Enhancement), a self-SR method that addresses each of these factors. ECLARE estimates the slice profile from the 2D-acquired multi-slice MR volume, trains a network to learn the mapping from low-resolution to high-resolution in-plane patches from the same volume, and performs SR with anti-aliasing. We compared ECLARE to cubic B-spline interpolation, SMORE, and other contemporary SR methods. We used realistic and representative simulations so that quantitative performance against a ground truth could be computed, and ECLARE outperformed all other methods in both signal recovery and downstream tasks. On real data for which there is no ground truth, ECLARE demonstrated qualitative superiority over other methods as well. Importantly, as ECLARE does not use external training data it cannot suffer from domain shift between training and testing. Our code is open-source and available at https://www.github.com/sremedios/eclare.
@@ 1 Introduction
Magnetic resonance (MR) imaging (MRI) is an essential clinical tool in diagnosis, monitoring, management and research in many neurological diseases. 2D MR neuroimaging uses slice selection to acquire 2D slices according to a specific slice profile (Brown et al., 2014). When multiple slices are acquired, these slices are stacked to create a 3D volume. In contrast, 3D MRI uses two phase encode directions to acquire 3D volumes directly (Elmaoğlu and Çelik, 2011). Although 3D MRI has many advantages, including greater utility in downstream processing, multi-slice 2D MRI still accounts for the majority of clinical scans because it 1) is faster to acquire; 2) offers clinically relevant tissue contrasts difficult to acquire in 3D; 3) has relaxed hardware requirements; 4) is very familiar to clinicians; and 5) is usually less sensitive to motion artifacts. In the national multi-site TREAT-MS pragmatic clinical trial on multiple sclerosis (MS) (Mowry and Newsome, 2024), for example, of 634 baseline scan sessions—all of which were ordered by different MS clinicians to verify MS diagnosis—only 58% of the sessions contain a 3D MR image and 81% of all reviewable scans obtained in these sessions are multi-slice 2D MRI. Multi-slice 2D MRI are also common in pre-clinical imaging studies (Stringer et al., 2021).
@@
Multi-slice 2D MR neuroimages can be strikingly anisotropic. Although in-plane resolutions for 2D MRI are frequently less than 1 mm in both phase-encode and readout directions, slice thicknesses may be 3–8 mm and many acquisitions additionally have slice gaps of 0.5–2 mm. Automated processing pipelines and algorithms, which are often designed and trained for isotropic image volumes, perform sub-optimally on such anisotropic volumes (even when interpolated to have isotropic voxel spacing). Super-resolution (SR) aims to solve this problem by estimating high-resolution (HR) images from low-resolution (LR) acquisitions. Most contemporary work on SR for MRI focuses on deep neural networks. Convolutional neural networks (CNNs) form the backbone of many approaches (Sood and Rusu, 2019; Du et al., 2020), including architectures such as generative adversarial networks (GANs) (Zhang et al., 2022; Chen et al., 2020; Lyu et al., 2020), normalizing flows (Lugmayr et al., 2020; Kim and Son, 2021; Liu et al., 2023; Ko et al., 2023), and denoising diffusion probabilistic models (DDPMs) (Wu et al., 2023; Wang et al., 2023; Wu et al., 2024a; Chang et al., 2024; Han and Huang, 2024; Wu et al., 2024b). The transformer architecture (Lu et al., 2022; Yang et al., 2020; Yan et al., 2021; Li et al., 2022; Feng et al., 2021) is also common, sometimes used alongside CNNs (Zhou et al., 2023; Forigua et al., 2022; Cheng et al., 2023; Fang et al., 2022).
@@
Although many SR approaches demonstrate remarkable performance, none at present satisfy all desirable properties for SR of multi-slice 2D MRI. Most methods, for example, do not consider the slice profile and true through-plane resolution of 2D MRI. Indeed, many approaches conflate slice separation with through-plane resolution, but they are different concepts and must be handled differently. For example, methods evaluated on simulations that use LR data that is only downsampled from 3D volumes (i.e., with interpolation) and without proper simulation of slice profile and slice separation would be more properly labeled as “inpainting” or “slice imputation” rather than super-resolution. As well, many methods in the literature have been trained for super-resolution of single integer factors such as 2 and 4 (Sood and Rusu, 2019; Zhang et al., 2018; Du et al., 2020; Lugmayr et al., 2020; Wu et al., 2023; Lu et al., 2022; Yang et al., 2020; Zhou et al., 2023). Not only do these methods need to be retrained for each integer factor that is encountered, they also use techniques like power-of-two upsampling or pixel shuffle (Shi et al., 2016) that do not allow for non-integer factors. Finally, many methods are trained on specific MR tissue contrasts and often (in neuroimaging) on skull-stripped data. Those methods would need to be retrained and possibly re-evaluated for alternative tissue contrasts and for non-skull-stripped data. At present, there is no algorithm that handles all of these concerns for an arbitrary multi-slice 2D MR image volume.
@@
To address these concerns, we propose ECLARE (Efficient Cross-planar Learning for Anisotropic Resolution Enhancement), a deep learning algorithm to super-resolve multi-slice 2D MR volumes along the through-plane direction. ECLARE builds on a succession of self-SR approaches (Jog et al., 2016; Zhao et al., 2018, 2019, 2020; Remedios et al., 2021, 2023), in which all training data is extracted from the LR image volume. However, ECLARE adds several crucial features to properly model slice acquisition, account for non-integer super-resolution factors, and incorporate efficient network design elements. In particular, we use ESPRESO (Han et al., 2021, 2023) to estimate the relative slice selection profile. We use this profile to model slice thickness and gap and to simulate training data from HR in-plane patches extracted from the anisotropic input volume. These data are used to train an efficient SR neural network architecture based on “wide activation SR” (WDSR) (Yu et al., 2018). The model is then applied to both canonical through-planes, yielding two super-resolved volumes that are then averaged to create the final output. Underlying all interpolation operations in our approach is a novel field-of-view (FOV)-aware interpolation algorithm that respects the FOV center and target sample spacing. The main contributions of this paper are the following:
@@
The sum of our contributions is a method that super-resolves an anisotropic volume without the need of external training data. This work extends a previously published extension to SMORE (Remedios et al., 2023).
@@ 2 Related Work
In this work, we consider SR for 2D multi-slice MRI. While there is a body of literature discussing SR approaches focusing on training a model with external data, we do not discuss it here. Such approaches create paired LR-HR data through simulations of large datasets, perform supervised learning with or without consideration of arbitrary scale factors, and either must be applied to a similar domain as the training set or undergo domain adaptation for practical use. Alternatively, in this paper we consider methods that aim to be applicable “out-of-the-box.” Accordingly, our discussion of related work is limited to self-supervised or foundational models for SR of 2D anisotropic MR volumes. Thus, at test-time, the existence of relevant external training data is not assumed.
2.1 Self-SR
@@
Self-SR in MRI was first explored (Jog et al., 2016) by fitting an anchored neighborhood regression (Timofte et al., 2013) model to learn the mapping from LR to HR patches, followed by Fourier burst accumulation (FBA) (Delbracio and Sapiro, 2015) to aggregate estimated high frequency Fourier coefficients into the super-resolved volume. The related work SMORE (Zhao et al., 2018, 2020) reinterpreted self-SR using CNNs, which greatly increased the output accuracy and visual appearance. SMORE trains two enhanced deep residual networks (EDSRs) (Lim et al., 2017): one for anti-aliasing and one for SR. These networks are applied in cascade to orthogonal through-plane slices, and all intermediate results are fused with FBA. Furthermore, although SMORE can be a self-SR method, Zhao et al. (2020) found that pre-training on similar data (i.e., human head MR images) and fine-tuning on the anisotropic volume at test-time allowed fewer training iterations and achieved better results.
2.2 INR
@@
Implicit neural representations (INR) (Park et al., 2019; Mildenhall et al., 2020; Sitzmann et al., 2020; Chen et al., 2021; Wu et al., 2022) learn the mapping from coordinate to intensity. Since coordinates can be sampled at any resolution, this has inspired several works related to arbitrary-scale SR (Tan et al., 2020; Zhu et al., 2021; Wu et al., 2022; Liu et al., 2024). Recently, multi-contrast INR (McGinnis et al., 2023) was proposed to super-resolve subjects with multiple available contrasts from the same session, ideally acquired at different orientations (e.g., sagittal $T_{1}$-w and coronal T2-w). However, that method also supports the use of a single contrast, training in a similar manner to self-SR methods. INR approaches cannot directly be used for SR, though. Since they train on bandlimited, sampled data, they cannot be expected to produce reasonable frequencies higher than the training data frequency cutoff, regardless of the queried sampling spacing. Furthermore, often the LR image exhibits blurring and aliasing, so the INR method also must correct for these factors rather than directly estimate the blurred, aliased intensity value.
2.3 SynthSR
@@
SynthSR (Iglesias et al., 2023) is a joint synthesis and SR model. Training data for SynthSR includes real MPRAGE head volumes and their corresponding parcellations. Domain randomization uses a Gaussian mixture model to produce synthetic MR-like images of various contrasts that are then augmented with a wide variety of simulated slice thicknesses, slice separations, bias fields, and additive noise. Through domain randomization, SynthSR maps an MR image of arbitrary contrast to a 1 mm isotropic T1-w MPRAGE scan.
@@ 3 Methods
Our proposed method ECLARE improves resolution along the through-plane axis of an anisotropic 2D-acquired multi-slice MR image volume. First, in Sec. 3.1 we model the relationship between HR and LR volumes, which is fully characterized by the slice selection profile and the slice separation. In Sec. 3.2, we describe how we approximate the slice selection profile and the mapping from HR to LR (the “forward model”). Then, in Sec. 3.3, we describe the different choices available for interpolation strategies and propose FOV-aware interpolation. We leverage a method named ESPRESO (Han et al., 2023), which produces a digital estimate of the slice selection profile. We then implement an approximation of our forward model with digital convolutions and our proposed FOV-aware interpolation strategy. In Sec. 3.4, we describe the method to simulate paired LR and HR 2D patches and perform supervised learning to train a 2D deep CNN. In Sec. 3.5, we show how ECLARE uses this 2D CNN to super-resolve the 3D volume along the two canonical through planes. Finally, in Sec. 3.6 we describe implementation details in the final ECLARE algorithm.
Figure 1: A flowchart of our proposed method. The anisotropic input volume is fed independently into each of the three steps. First, in (A) (Sec. 3.2), we estimate the slice excitation profile with ESPRESO. Second, in (B) (Sec. 3.4), we extract HR in-plane 2D patches and use the PSF estimated from (A) to create paired training data. This training data is used to train the network with supervised learning. Third, in (C) (Sec. 3.5), we extract LR through-plane 2D slices and super-resolve them with the trained network from (B).The super-resolved slices are stacked and averaged, yielding the super-resolved output volume.
@@ 3.1 Modeling the imaging process
A 2D-acquired MR image volume is a stack of independent 2D slices. In-plane, data is acquired in k-space and transformed to the image domain through the inverse Fourier transform. For through-plane, the relationship between slices is entirely characterized in the image domain. Along the through-plane axis, a magnetic gradient causes atomic nuclei to resonate at different frequencies. This means that resonant frequencies along this axis correspond to different spatial locations. Slice selection is a bandpass filter that selects frequencies (again, corresponding to spatial location) that are to be excited. The slice selection profile, (equivalently, “slice profile”) is the shape of the bandpass filter. Although the ideal form of the slice selection profile is a rect function, which would excite only the through-plane positions corresponding to the slice of interest and nothing else, in practice this is unachievable. Instead, common slice profiles take the shape of truncated Gaussians or use the Shinnar-Le Roux algorithm (Pauly et al., 1991; Shinnar et al., 1989) to approximate a rect. Assuming homogeneity in the imaging process, within its frequency band the slice selection profile excites all atomic nuclei equally. This means all in-plane measurements are subject to the same slice thickness and location.
@@
The desired image volume $x$ is HR and isotropic with resolution $a\times a\times a$ mm. In this work we consider it to also be a 2D-acquired multi-slice MR volume with slice thickness equal to the slice separation (i.e., no gap). In reality, acquiring such image volumes is not practical; 3D acquisitions are preferred. However, for simplification we consider only 2D-acquisitions. Now, the observed image volume $y$ is LR and anisotropic. Its resolution $a\times a\times b$ mm is the same as the HR volume in-plane but worse along the through-plane axis: $b>a$. Additionally, $y$ may have slice gap, so the slice separation may be greater than the slice thickness. We model 2D multi-slice acquisitions as a discrete convolution between the unobserved HR image volume and the slice selection profile:
@@
where $\downarrow_{r}$ is the subsampling ratio of $r$. The full-width-at-half-max (FWHM) of $h$ determines the slice thickness (alongside the HR slice thickness $a$) and $r$ is a ratio between the HR slice separation $s^{\ast}$ and the LR slice separation $s$. We define this factor as $r=\frac{s}{s^{\ast}}$. If $h$ has a smaller FWHM than the separation $s$, then Eq. 1 models slice gap. Similarly, a FWHM of $h$ equal to or greater than $s$ respectively models adjacent slices and slice overlap.

### v2

@@ Abstract
In clinical imaging, magnetic resonance (MR) image volumes are often acquired as stacks of 2D slices with decreased scan times, improved signal-to-noise ratio, and image contrasts unique to 2D MR pulse sequences. While this is sufficient for clinical evaluation, automated algorithms designed for 3D analysis perform poorly on multi-slice 2D MR volumes, especially those with thick slices and gaps between slices. Super-resolution (SR) methods aim to address this problem, but previous methods do not address all of the following: slice profile shape estimation, slice gap, domain shift, and non-integer or arbitrary upsampling factors. In this paper, we propose ECLARE (Efficient Cross-planar Learning for Anisotropic Resolution Enhancement), a self-SR method that addresses each of these factors. ECLARE uses a slice profile estimated from the multi-slice 2D MR volume, trains a network to learn the mapping from low-resolution to high-resolution in-plane patches from the same volume, and performs SR with anti-aliasing. We compared ECLARE to cubic B-spline interpolation, SMORE, and other contemporary SR methods. We used realistic and representative simulations so that quantitative performance against ground truth can be computed, and ECLARE outperformed all other methods in both signal recovery and downstream tasks. Importantly, as ECLARE does not use external training data it cannot suffer from domain shift between training and testing. Our code is open-source and available at https://www.github.com/sremedios/eclare.
@@ 1 Introduction
Magnetic resonance (MR) imaging (MRI) is an essential clinical tool in diagnosis, monitoring, management, and research across many health conditions. Multi-slice 2D MRI uses slice selection [1] to acquire multiple evenly spaced 2D slices, which are stacked to create a 3D volume. In contrast, 3D MRI uses two phase encode directions to acquire 3D volumes directly [2]. Although 3D MRI has many advantages, including greater utility in downstream processing, multi-slice 2D MRI still accounts for the majority of clinical scans because it: 1) is faster to acquire; 2) offers clinically relevant tissue contrasts difficult to acquire in 3D [3]; 3) has relaxed hardware requirements; 4) is very familiar to clinicians; and 5) can be less sensitive to motion artifacts. For these reasons, multi-slice 2D MRI are common in pragmatic clinical trials [4] and in pre-clinical imaging studies [5].
@@
Multi-slice 2D MR neuroimages can be strikingly anisotropic. Although in-plane resolutions for 2D MRI are frequently less than 1 mm in both phase-encode and readout directions, slice thicknesses may be 3–8 mm and many acquisitions additionally have slice gaps of 0.5–2 mm. Automated processing pipelines, which are often designed and trained for isotropic image volumes, perform sub-optimally on such anisotropic volumes (even when interpolated to isotropic voxel spacing). Super-resolution (SR) aims to solve this problem by estimating high-resolution (HR) images from low-resolution (LR) acquisitions. Most contemporary work on SR for MRI focuses on deep neural networks. Convolutional neural networks (CNNs) form the backbone of many approaches [6, 7], including generative adversarial networks (GANs) [8, 9, 10], normalizing flows [11, 12, 13, 14], and denoising diffusion probabilistic models (DDPMs) [15, 16, 17, 18, 19, 20]. The transformer architecture [21, 22, 23, 24, 25] is also common, sometimes used alongside CNNs [26, 27, 28, 29].
Figure 1: A flowchart of our proposed method. The anisotropic input volume is fed independently into each of the three steps. First, in (A) (Sec. 3.2), we estimate the slice excitation profile with ESPRESO [30]. Second, in (B) (Sec. 3.4), we extract HR in-plane 2D patches and use the PSF estimated from (A) to create paired training data. This training data is used to train the network $f_{\theta}$ with supervised learning. Third, in (C) (Sec. 3.5), we extract LR through-plane 2D slices and super-resolve them with the trained network $f_{\theta}$ from (B).The super-resolved slices are stacked and averaged, yielding the super-resolved output volume.
@@
Although many SR approaches demonstrate remarkable performance, none at present satisfy all desirable properties for SR of multi-slice 2D MRI. Most methods, for example, do not consider the slice profile and true through-plane resolution of 2D MRI. Indeed, many approaches conflate slice separation with through-plane resolution, but they are different concepts and must be handled differently. For example, methods evaluated on simulations that use LR data that is only downsampled from 3D volumes (i.e., with interpolation) and without proper simulation of slice profile and slice separation would be more properly labeled as “inpainting” or “slice imputation” rather than super-resolution. As well, many methods in the literature have been trained for super-resolution of single integer factors such as 2 and 4 [6, 31, 11, 15, 21, 26]. Not only do these methods require retraining for each integer factor that is encountered, they also use techniques like power-of-two upsampling or pixel shuffle [32] that do not allow for non-integer factors. Finally, many methods are trained on specific MR tissue contrasts and often (in neuroimaging) on skull-stripped data, necessitating retraining and possibly re-evaluation for alternative tissue contrasts and for with-skull data. At present, there is no algorithm that handles all of these concerns for an arbitrary multi-slice 2D MR image volume.
@@
To address these concerns, we propose ECLARE (Efficient Cross-planar Learning for Anisotropic Resolution Enhancement), a deep learning algorithm to super-resolve multi-slice 2D MR volumes along the through-plane direction. ECLARE builds on a succession of self-SR approaches [33, 34, 35, 36], in which all training data is extracted from the HR in-plane slices within a LR image volume. However, ECLARE adds several crucial features to properly model slice acquisition, account for non-integer super-resolution factors, and incorporate efficient network design elements. In particular, we use ESPRESO [30], a GAN-based algorithm previously developed by our team, to estimate the relative slice selection profile. We use this profile to model slice thickness and gap and to simulate training data from HR in-plane patches extracted from the anisotropic input volume. These data are used to train an efficient SR neural network architecture based on “wide activation SR” (WDSR) [37]. The model is then applied to both canonical through-planes, yielding two super-resolved volumes that are then averaged to create the final output. Underlying all interpolation operations in our approach is a novel field-of-view (FOV)-aware interpolation algorithm that respects the image coordinate system and target sample spacing. The main contributions of this paper are the following:
@@
Our proposed method super-resolves an anisotropic volume without external training data and can serve as an alternative to spatial interpolation with improved quality.
@@ 2 Related Work
In this work, we consider SR for 2D multi-slice MRI. While there is a body of literature discussing SR approaches focusing on training a model with external data, we do not discuss it here. Such approaches create paired LR-HR data through simulation from large datasets, perform supervised learning with or without consideration of arbitrary scale factors, and either must be applied to a similar domain as the training set or undergo domain adaptation for practical use. Alternatively, in this paper we consider methods that aim to be applicable “out-of-the-box.” Accordingly, our discussion of related work is limited to self-supervised or foundational models for SR of 2D anisotropic MR volumes. Thus, at test-time, the existence of relevant external training data is not assumed.
Figure 2: A 1D demonstration of the differences between interpolation implementations. The input has 6 pixels and its FOV center is located at 2.5, indicated by the red dashed line. Its sample spacing is 1, indicated by the purple arrow. Its extents occur at -0.5 and 5.5, indicated by the blue dashed line. In this demonstration, the desired sample spacing for upsampling is 0.7, and the desired sample spacing for downsampling is 1.429. The implementation scipy.zoom [38] maintains the FOV center and boundary sample locations, but changes the sample spacing to 0.625 and 1.667 for upsampling and downsampling, respectively. The implementation torch.nn.functional.interpolate [39] anchors the FOV extent to start at 0.5, shifting the center from the desired FOV center. Our proposed method (Sec. 3.3) maintains both the sample spacing and the FOV center.
@@
Self-SR in MRI was first explored [33] by fitting an anchored neighborhood regression [40] model to learn the mapping from LR to HR patches, followed by Fourier burst accumulation (FBA) [41] to aggregate estimated high frequency Fourier coefficients into the super-resolved volume. The related work SMORE [34] reinterpreted self-SR using CNNs, which greatly increased the output accuracy and visual appearance. SMORE trains two enhanced deep residual networks (EDSRs) [42]: one for anti-aliasing and one for SR. These networks are applied in cascade to orthogonal through-plane slices, and all intermediate results are fused with FBA. Furthermore, although SMORE can be a self-SR method, [34] found that pre-training on similar data (i.e., human head MR images) and fine-tuning on the anisotropic volume at test-time allowed fewer training iterations and achieved better results.
@@
Implicit neural representations (INR) [43, 44, 45] learn the mapping from coordinate to intensity. Since coordinates can be sampled at any resolution, this has inspired several works related to arbitrary-scale SR [46, 45, 47]. Recently, multi-contrast INR [48] was proposed to super-resolve subjects with multiple available contrasts from the same session, ideally acquired at different orientations (e.g., sagittal $T_{1}$-w and coronal T2-w). However, that method also supports the use of a single contrast, training in a similar manner to self-SR methods.
@@
SynthSR [49] is a joint synthesis and SR model. Training data for SynthSR includes real MPRAGE head volumes and their corresponding parcellations. Domain randomization uses a Gaussian mixture model to produce synthetic MR-like images of various contrasts alongside augmentations such as simulated anisotropy, slice separations, bias fields, and additive noise. Through domain randomization, SynthSR maps an MR image of arbitrary contrast to a 1 mm isotropic T1-w MPRAGE scan.
@@ 3 Methods
Our proposed ECLARE improves resolution along the through-plane axis of an anisotropic 2D-acquired multi-slice MR image volume. To explain ECLARE, we first model the relationship between HR and LR volumes (Sec. 3.1). We then describe how to approximate the slice selection profile (Sec. 3.2) and FOV-aware interpolation (Sec. 3.3). We approximate the slice selection profile with ESPRESO [30] and use it to train a CNN with in-plane HR patches and their simulated LR counterparts (Sec. 3.4). Finally, we show how ECLARE uses this 2D CNN to super-resolve the 3D volume (Sec. 3.5) and describe implementation details (Sec. 3.6).
Figure 3: A diagram of our proposed architecture, based on WDSR [37]. The low resolution image $y$ is first processed by a convolution to produce a feature map with 256 channels. It then passes through several WDSR blocks. A final convolution is done to produce the correct number of channels for pixel shuffle. Finally, the result is downsampled and added to an upsampled $y$ to yield the output. In our implementation, we use 16 WDSR blocks.
@@ 3.1 Modeling the imaging process
A 2D-acquired MR image volume is a stack of independent 2D slices. In-plane, data is acquired in k-space and transformed to the image domain through the inverse Fourier transform. For through-plane, the relationship between slices is entirely characterized in the image domain. Common slice profiles take the shape of truncated Gaussians or use the Shinnar-Le Roux algorithm [50, 51] to approximate a rect. All in-plane measurements are subject to the same slice thickness and location.
@@
An ideal multi-slice 2D MR volume $x$ would be acquired in an isotropic fashion with resolution $a\times a\times a$ mm. For practical reasons (time and physical constraints), however, a real multi-slice 2D MR volume $y$ is acquired with thicker slices—i.e., $a\times a\times b$ mm where $b>a$. A further problem is that $y$ may have slice gaps where the slice separation in the through-plane direction is greater than $b$, the slice thickness. We model 2D multi-slice acquisitions as a discrete convolution between the unobserved HR image volume and the slice selection profile:
@@
where $\downarrow_{r}$ is the subsampling ratio of $r$. The full-width-at-half-max (FWHM) of $h$ determines the slice thickness and $r$ is a ratio between the HR slice separation $s^{\ast}$ and the LR slice separation $s$, defined as $r=\frac{s}{s^{\ast}}$. If $h$ has a smaller FWHM than the separation $s$, then Eq. 1 models slice gap. Similarly, a FWHM of $h$ equal to or greater than $s$ respectively models adjacent slices and slice overlap.